A beam of an unknown, brownish material is 1.0 m long at a temperature of 35 degrees C. On a hot day, the temperature rises to 45 degrees C. The beam's length is observed to increase by 2.1 mm. What is the material of the beam?
Answer #1

l=lo (It QDT) l-lo falo &DT 2.1 - 1x2 (45-35) 12 = 2.1810-toc- Check from the beovided table, which material has a thenal Cof